Matt is twice as old as his daughter Sally. Ten years ago, the sum of their ages was 31. How old is each person now?
Answer by Cromlix(4381)

Hi there,
Matt is twice as old as his daughter Sally.
M = 2S
Ten years ago, the sum of their ages was 31.
M - 10 + S - 10 = 31
Replace M with 2S from top equation
2S - 10 + S - 10 = 31
Collect like terms
3S = 31 + 20
3S = 51
S = 17
Matt is 34 years old
Sally is 17 years old
Hope this helps :-)